# itemsep=1pt,topsep=2pt,parsep=1pt,partopsep=1pt

### v2

@@ 8.2. Second moment of $u^{\delta}[f]$
(8.10) $\delta^{2}\iint_{{\mathbbm{R}}^{2}}f(\delta x)f(\delta y)K^{\delta}(x,y){\mathrm{d}}x{\mathrm{d}}y,$
with $x_{0}:=x$, $y_{0}:=y$, and
(8.11) $K^{\delta}(x,y):=\sum\limits_{k=1}^{\infty}\beta_{\delta}^{2k}\idotsint\limits_{0<t_{1}<\cdots<t_{k}<\delta^{-2}\atop x_{1},\cdots,x_{k},y_{1},\cdots,y_{k}\in{\mathbbm{R}}}\prod\limits_{i=1}^{k}g_{t_{i}-t_{i-1}}(x_{i}-x_{i-1})g_{t_{i}-t_{i-1}}(y_{i}-y_{i-1})\rho(x_{i})\rho(y_{i}){\mathrm{d}}x_{i}{\mathrm{d}}y_{i}{\mathrm{d}}t_{i}.$
By a change of variable $t_{1}=\delta^{-2}t$ and the translation invariance of $g_{t}(x)$, we have that
(8.12) $K^{\delta}(x,y)=\int_{0}^{1}{\mathrm{d}}t\int_{{\mathbbm{R}}^{2}}g_{t}(\delta(x^{\prime}-x))g_{t}(\delta(y^{\prime}-y))\rho(x^{\prime})\rho(y^{\prime})K_{1-t}^{\delta}(x^{\prime},y^{\prime}){\mathrm{d}}x^{\prime}{\mathrm{d}}y^{\prime},$
with $x_{0}:=x^{\prime}$, $y_{0}:=y^{\prime}$, and
(8.13) $K^{\delta}_{1-t}(x^{\prime},y^{\prime}):=\sum\limits_{k=0}^{\infty}\beta_{\delta}^{2(k+1)}\!\!\!\!\idotsint\limits_{0<t_{1}<\cdots<t_{k}<\delta^{-2}(1-t)\atop x_{1},\cdots,x_{k},y_{1},\cdots,y_{k}\in{\mathbbm{R}}}\prod\limits_{i=1}^{k}g_{t_{i}-t_{i-1}}(x_{i}-x_{i-1})g_{t_{i}-t_{i-1}}(y_{i}-y_{i-1})\rho(x_{i})\rho(y_{i}){\mathrm{d}}x_{i}{\mathrm{d}}y_{i}{\mathrm{d}}t_{i},$
and the integral is set to be 00 when $k=0$. We will see that (8.13) plays a the role of $\overline{U}_{N}(n)$ (see (2.24)) in the discrete setting with $N=\lfloor\delta^{-2}\rfloor$ and $n=\lfloor t\rfloor$.
Now write
$\bar{r}(t)=\sup_{x^{\prime},y^{\prime}\in\text{supp}(\rho)}\iint_{{\mathbbm{R}}^{2}}g_{t}(x-x^{\prime})g_{t}(y-y^{\prime})\rho(x)\rho(y){\mathrm{d}}x{\mathrm{d}}y=\frac{1}{2\pi t}(1+o(1)),\quad\text{and}~{}\bar{{\mathcal{R}}}_{\delta}=\int_{0}^{\delta^{-2}}\bar{r}(t){\mathrm{d}}t,$
and note that $\beta_{\delta}^{2}\bar{{\mathcal{R}}}_{\delta}\leq 1+\frac{c}{\log\delta^{-2}}$, since $\beta_{\delta}^{2}=\frac{\pi}{\log\delta^{-1}}+\frac{\theta+o(1)}{(\log\delta^{-1})^{2}}$. Furthermore, by (8.7)-(8.9), for $t,x^{\prime},y^{\prime}$ uniformly, we can restrict the summation in (8.13) to $k\leq A\log\delta^{-1}$ for $A$ sufficiently large and restrict the integral there for $t_{1},\cdots,t_{k}$ to
(8.14) $\begin{split}{\mathcal{J}}^{\delta,k}_{1-t}:=\{0<t_{1}<\cdots<t_{k}<&\delta^{-2}(1-t):t_{1}>(\log\delta^{-1})^{2},\text{and}~{}\forall~{}1\leq i\leq r-1,\\ &\text{either}~{}t_{i}-t_{i-1}>(\log\delta^{-1})^{2}~{}\text{or}~{}t_{i+1}-t_{i}>(\log\delta^{-1})^{2}\}.\end{split}$
We denote (8.13) by $K_{1-t,{\mathcal{J}}}^{\delta,A}(x^{\prime},y^{\prime})$ after the summation and integral are restricted. Let
(8.15) $r(t):=\idotsint_{{\mathbbm{R}}^{4}}\rho(x^{\prime})\rho(y^{\prime})g_{t}(x-x^{\prime})g_{t}(y-y^{\prime})\rho(x)\rho(y){\mathrm{d}}x{\mathrm{d}}x^{\prime}{\mathrm{d}}y{\mathrm{d}}y^{\prime}=\frac{1}{2\pi t}(1+o(1)).$
@@
We then treat the integral in $K_{1-t,{\mathcal{J}}}^{\delta,A}$ in the following manners:
@@
(8.16) $\begin{split}\widehat{K}_{1-t}^{\delta}&=\sum\limits_{k=0}^{\infty}\beta_{\delta}^{2(k+1)}\iint\limits_{0<t_{1}<\cdots<t_{k}<\delta^{-2}(1-t)}\prod\limits_{i=1}^{k}r(t_{i}-t_{i-1}){\mathrm{d}}t_{i}\\ &=\frac{2\pi+o(1)}{\log\delta^{-2}}\sum\limits_{k=0}^{\infty}(\beta_{\delta}^{2}\bar{\mathcal{R}}_{\delta})^{k}{\mathbf{P}}({\mathcal{T}}_{1}^{\delta}+\cdots+{\mathcal{T}}_{k}^{\delta}\leq\delta^{-2}(1-t)),\end{split}$
@@
(8.17) $\lim\limits_{\delta\to 0}\widehat{K}_{1-t}^{\delta}=2\pi\int_{0}^{\infty}e^{\vartheta u}{\mathbf{P}}(Y_{u}\leq 1-t){\mathrm{d}}u=2\pi\int_{0}^{\infty}e^{\vartheta u}{\mathrm{d}}u\int_{0}^{1-t}\alpha_{u}(r){\mathrm{d}}r=2\pi\int_{0}^{1-t}G_{\vartheta}(r){\mathrm{d}}r.$
Hence, by a change of variables $\delta x=x_{0}$, $\delta y=y_{0}$ and $r=s-t$, the second moment of $u^{\delta}[f]$ equals to $(\int_{\mathbbm{R}}f(x){\mathrm{d}}x)^{2}$ plus $(1+o(1))2\pi$ multiplying
(8.18) ∫⋯∫$\begin{split}&\idotsint_{{\mathbbm{R}}^{4}}f(x_{0})f(y_{0})g_{t}(\delta x^{\prime}-x_{0})g_{t}(\delta y^{\prime}-y_{0})\rho(x^{\prime})\rho(y^{\prime}){\mathrm{d}}x_{0}{\mathrm{d}}y_{0}{\mathrm{d}}x^{\prime}{\mathrm{d}}y^{\prime}\iint\limits_{0<t<s<1}G_{\vartheta}(s-t){\mathrm{d}}s{\mathrm{d}}t\\ \longrightarrow&\idotsint\limits_{0<t<s<1\atop x,y\in{\mathbbm{R}}}f(x)f(y)g_{t}(x)g_{t}(y)G_{\vartheta}(s-t){\mathrm{d}}s{\mathrm{d}}t{\mathrm{d}}x{\mathrm{d}}y,\quad\text{as}~{}\delta\to 0.\end{split}$
We now determine $\vartheta$ and then the second moment computations will be completed. Note that
(8.19) ${\mathcal{R}}_{\delta}=\idotsint_{{\mathbbm{R}}^{4}}\rho(x^{\prime})\rho(y^{\prime})\rho(x)\rho(y){\mathrm{d}}x^{\prime}{\mathrm{d}}y^{\prime}{\mathrm{d}}x{\mathrm{d}}y\int_{0}^{\delta^{-2}}g_{t}(x-x^{\prime})g_{t}(y-y^{\prime}){\mathrm{d}}t.$
We have that by the end of [CSZ19b, Section 8.2],
@@
$R_{\delta}=\frac{1}{2\pi}\Big{(}\log\delta^{-2}+\log 2-\gamma+\idotsint_{{\mathbbm{R}}^{4}}\!\rho(x^{\prime})\rho(y^{\prime})\log\!\Big{[}\frac{1}{(x-x^{\prime})^{2}+(y-y^{\prime})^{2}}\Big{]}\rho(x)\rho(y){\mathrm{d}}x^{\prime}{\mathrm{d}}y^{\prime}{\mathrm{d}}x{\mathrm{d}}y+o(1)\Big{)}.$
@@
(8.20) $\vartheta=\log 2-\gamma+\idotsint_{{\mathbbm{R}}^{4}}\rho(x^{\prime})\rho(y^{\prime})\log\Big{[}\frac{1}{(x-x^{\prime})^{2}+(y-y^{\prime})^{2}}\Big{]}\rho(x)\rho(y){\mathrm{d}}x^{\prime}{\mathrm{d}}y^{\prime}{\mathrm{d}}x{\mathrm{d}}y+\frac{\theta}{2\pi}.$
8.3. Approximation by coarse-grained model
@@
We partition the time interval $[0,\delta^{-2}]$ into mesoscopic intervals
(8.21) ${\mathcal{I}}_{\delta,\varepsilon}(\mathrm{i}):=[(\mathrm{i}-1)\varepsilon\delta^{-2},\varepsilon\delta^{-2}],\quad\mathrm{i}=1,\cdots,\frac{1}{\varepsilon}.$
@@
(8.22) $\begin{split}u^{\delta}[f]=&\int_{\mathbbm{R}}f(x){\mathrm{d}}x+\delta\sum\limits_{k=1}^{1/\varepsilon}\sum\limits_{1\leq\mathrm{i}_{1}<\cdots<\mathrm{i}_{k}<\frac{1}{\varepsilon}}\sum\limits_{j_{0}:=0<j_{1}<\cdots<j_{k}}\int_{{\mathbbm{R}}}f(\delta x_{0}){\mathrm{d}}x_{0}\times\\ &\bigg{(}\prod\limits_{\ell=1}^{k}\beta_{\delta}^{j_{\ell}-j_{\ell-1}}\idotsint\limits_{\quad t_{j_{\ell-1}+1}<\cdots<t_{j_{\ell}}\in{\mathcal{I}}_{\delta,\varepsilon}(\mathrm{i}_{2})\atop x_{j_{\ell-1}+1},\cdots,x_{j_{\ell}}\in{\mathbbm{R}}}\prod\limits_{i=j_{\ell-1}+1}^{j_{\ell}}g_{t_{i}-t_{i-1}}(x_{i}-x_{i-1})\rho(x_{i}){\mathrm{d}}x_{i}{\mathrm{d}}W_{t_{i}}\bigg{)}.\end{split}$
By the independent increments of $W_{t}$, when computing the second moment of (8.22), the cross terms related to different $\mathrm{i}_{1}<\cdots<\mathrm{i}_{k}$ and $\mathrm{i}^{\prime}_{1}<\cdots<\mathrm{i}^{\prime}_{l}$ have no contributions.By the second moment computations in Subsection 8.2, we have an continuous analogue of Lemma 5.2, that is, as $\delta\to 0$, for fixed ${\mathbf{I}}=(\mathrm{i}_{1},\cdots,\mathrm{i}_{k})$, the related second moment converges to ${\mathcal{I}}_{\varepsilon}^{f}({\mathbf{I}})$ defined in (5.5). Thus, we can restrict the summation for $\mathrm{i}_{1},\cdots,\mathrm{i}_{k}$ to $\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}$ defined in (2.42) as we have done in Section 5.
@@
(8.23) $\Theta(\vec{\mathrm{i}}):=\begin{cases}\displaystyle\frac{\delta}{\sqrt{\varepsilon}}\sum\limits_{k=2}^{\infty}\beta_{\delta}^{k}\int_{{\mathbbm{R}}}\rho(x_{1}){\mathrm{d}}x_{1}\idotsint\limits_{t_{1}<\cdots<t_{k}\in{\mathcal{I}}_{\delta,\varepsilon}(\mathrm{i})\atop x_{2},\cdots,x_{k}\in{\mathbbm{R}}}\prod\limits_{i=2}^{k}g_{t_{i}-t_{i-1}}(x_{i}-x_{i-1})\rho(x_{i}){\mathrm{d}}x_{i}{\mathrm{d}}W_{t_{i}},&~{}~{}\text{if}~{}|\vec{\mathrm{i}}|=1,\\ \displaystyle\begin{split}\frac{\delta}{\sqrt{\varepsilon}}\sum\limits_{k=1}^{\infty}\beta_{\delta}^{k}&\int_{{\mathbbm{R}}}\rho(x_{1}){\mathrm{d}}x_{1}\idotsint\limits_{t_{1}<\cdots<t_{k}\in{\mathcal{I}}_{\delta,\varepsilon}(\mathrm{i})\atop x_{2},\cdots,x_{k}\in{\mathbbm{R}}}\prod\limits_{i=1}^{k}g_{t_{i}-t_{i-1}}(x_{i}-x_{i-1})\rho(x_{i}){\mathrm{d}}x_{i}{\mathrm{d}}W_{t_{i}}\times\\ &\sum\limits_{j=1}^{\infty}\beta_{\delta}^{j}\int_{{\mathcal{I}}_{\delta,\varepsilon}(\mathrm{i}^{\prime})}\int_{\mathbbm{R}}g_{s_{1}-t_{k}}(y_{1}-x_{k})\rho(y_{1}){\mathrm{d}}y_{1}{\mathrm{d}}W_{s_{1}}\times\\ &\idotsint\limits_{s_{2}<\cdots<s_{j}\in{\mathcal{I}}_{\delta,\varepsilon}(\mathrm{i}^{\prime})\atop y_{2},\cdots,y_{j}\in{\mathbbm{R}}}\prod\limits_{i=2}^{j}g_{s_{i}-s_{i-1}}(y_{i}-y_{i-1})\rho(y_{i}){\mathrm{d}}y_{i}{\mathrm{d}}W_{s_{i}}\end{split},&\text{if}~{}|\vec{\mathrm{i}}|>1,\end{cases}$

### v3

@@ 8.2 Second moment of $u^{\delta}[f]$
$\delta^{2}\iint_{{\mathbbm{R}}^{2}}f(\delta x)f(\delta y)K^{\delta}(x,y){\mathrm{d}}x{\mathrm{d}}y,$ (8.10)
where
$\begin{split}&K^{\delta}(x,y):=\\ \sum\limits_{k=1}^{\infty}\beta_{\delta}^{2k}\idotsint\limits_{0<t_{1}<\cdots<t_{k}<\delta^{-2}\atop x_{1},\cdots,x_{k},y_{1},\cdots,y_{k}\in{\mathbbm{R}}}&\prod\limits_{i=1}^{k}g_{t_{i}-t_{i-1}}(x_{i}-x_{i-1})g_{t_{i}-t_{i-1}}(y_{i}-y_{i-1})\rho(x_{i})\rho(y_{i}){\mathrm{d}}x_{i}{\mathrm{d}}y_{i}{\mathrm{d}}t_{i}\end{split}$ (8.11)
with the convention $x_{0}:=x$, $y_{0}:=y$. By a change of variable $t_{1}=\delta^{-2}t$ and the translation invariance of $g_{t}(x)$, we have that
$K^{\delta}(x,y)=\int_{0}^{1}{\mathrm{d}}t\int_{{\mathbbm{R}}^{2}}g_{t}(\delta(x^{\prime}-x))g_{t}(\delta(y^{\prime}-y))\rho(x^{\prime})\rho(y^{\prime})K_{1-t}^{\delta}(x^{\prime},y^{\prime}){\mathrm{d}}x^{\prime}{\mathrm{d}}y^{\prime},$ (8.12)
where
$\begin{split}&K^{\delta}_{1-t}(x^{\prime},y^{\prime}):=\beta_{\delta}^{2}+\sum\limits_{k=1}^{\infty}\beta_{\delta}^{2(k+1)}\\ \times\idotsint\limits_{0<t_{1}<\cdots<t_{k}<\delta^{-2}(1-t)\atop x_{1},\cdots,x_{k},y_{1},\cdots,y_{k}\in{\mathbbm{R}}}&\prod\limits_{i=1}^{k}g_{t_{i}-t_{i-1}}(x_{i}-x_{i-1})g_{t_{i}-t_{i-1}}(y_{i}-y_{i-1})\rho(x_{i})\rho(y_{i}){\mathrm{d}}x_{i}{\mathrm{d}}y_{i}{\mathrm{d}}t_{i}\end{split}$ (8.13)
with the convention $x_{0}:=x^{\prime}$, $y_{0}:=y^{\prime}$. We will see that (8.13) plays the role of $\overline{U}_{N}(n)$ (see (2.24)) in the discrete setting with $N=\lfloor\delta^{-2}\rfloor$ and $n=\lfloor t\rfloor$.
Now with $\bar{{\mathcal{R}}}_{\delta}=\int_{0}^{\delta^{-2}}\bar{r}(t){\mathrm{d}}t$, we write
$\bar{r}(t)=\sup_{x^{\prime},y^{\prime}\in\text{supp}(\rho)}\iint_{{\mathbbm{R}}^{2}}g_{t}(x-x^{\prime})g_{t}(y-y^{\prime})\rho(x)\rho(y){\mathrm{d}}x{\mathrm{d}}y=\frac{1}{2\pi t}(1+o(1)),$
and note that $\beta_{\delta}^{2}\bar{{\mathcal{R}}}_{\delta}\leq 1+\frac{c}{\log\delta^{-2}}$, since $\beta_{\delta}^{2}=\frac{\pi}{\log\delta^{-1}}+\frac{\theta+o(1)}{(\log\delta^{-1})^{2}}$. Furthermore, by (8.7)-(8.9), for $t,x^{\prime},y^{\prime}$ uniformly, we can restrict the summation in (8.13) to $k\leq A\log\delta^{-1}$ for $A$ sufficiently large and restrict the domain of integrals there to
$\begin{split}{\mathcal{J}}^{\delta,k}_{1-t}:=\{0<t_{1}&<\cdots<t_{k}<\delta^{-2}(1-t):t_{1}>(\log\delta^{-1})^{2},\text{and}~\forall~1\leq i\leq r-1,\\ &\text{either}~t_{i}-t_{i-1}>(\log\delta^{-1})^{2}~\text{or}~t_{i+1}-t_{i}>(\log\delta^{-1})^{2}\}.\end{split}$ (8.14)
We denote (8.13) by $K_{1-t,{\mathcal{J}}}^{\delta,A}(x^{\prime},y^{\prime})$ after the summation and integrals are restricted. Let
$r(t):=\idotsint_{{\mathbbm{R}}^{4}}\rho(x^{\prime})\rho(y^{\prime})g_{t}(x-x^{\prime})g_{t}(y-y^{\prime})\rho(x)\rho(y){\mathrm{d}}x{\mathrm{d}}x^{\prime}{\mathrm{d}}y{\mathrm{d}}y^{\prime}=\frac{1}{2\pi t}(1+o(1)).$ (8.15)
@@
We then treat the integrals in $K_{1-t,{\mathcal{J}}}^{\delta,A}$ in the following manners:
@@
$\begin{split}\widehat{K}_{1-t}^{\delta}&=\sum\limits_{k=0}^{\infty}\beta_{\delta}^{2(k+1)}\idotsint\limits_{0<t_{1}<\cdots<t_{k}<\delta^{-2}(1-t)}\prod\limits_{i=1}^{k}r(t_{i}-t_{i-1}){\mathrm{d}}t_{i}\\ &=\frac{2\pi+o(1)}{\log\delta^{-2}}\sum\limits_{k=0}^{\infty}(\beta_{\delta}^{2}{\mathcal{R}}_{\delta})^{k}{\mathbf{P}}({\mathcal{T}}_{1}^{\delta}+\cdots+{\mathcal{T}}_{k}^{\delta}\leq\delta^{-2}(1-t)),\end{split}$ (8.16)
@@
$\lim\limits_{\delta\to 0}\widehat{K}_{1-t}^{\delta}\!=\!2\pi\int_{0}^{\infty}\!\!e^{\vartheta u}{\mathbf{P}}(Y_{u}\leq 1\!-\!t){\mathrm{d}}u\!=\!2\pi\int_{0}^{\infty}\!\!e^{\vartheta u}{\mathrm{d}}u\int_{0}^{1-t}\!\!f_{u}(r){\mathrm{d}}r=2\pi\int_{0}^{1-t}\!\!G_{\vartheta}(r){\mathrm{d}}r.$ (8.17)
Hence, by a change of variables $\delta x=x_{0}$, $\delta y=y_{0}$ and r=str=s-t, the second moment of $u^{\delta}[f]$ equals to $(\int_{\mathbbm{R}}f(x){\mathrm{d}}x)^{2}$ plus $(1+o(1))2\pi$ multiplying by
$\begin{split}\idotsint_{{\mathbbm{R}}^{4}}&f(x_{0})f(y_{0})g_{t}(\delta x^{\prime}-x_{0})g_{t}(\delta y^{\prime}-y_{0})\rho(x^{\prime})\rho(y^{\prime}){\mathrm{d}}x_{0}{\mathrm{d}}y_{0}{\mathrm{d}}x^{\prime}{\mathrm{d}}y^{\prime}\!\!\!\!\iint\limits_{0<t<s<1}\!\!G_{\vartheta}(s-t){\mathrm{d}}s{\mathrm{d}}t\\ &\longrightarrow\idotsint\limits_{0<t<s<1\atop x,y\in{\mathbbm{R}}}f(x)f(y)g_{t}(x)g_{t}(y)G_{\vartheta}(s-t){\mathrm{d}}s{\mathrm{d}}t{\mathrm{d}}x{\mathrm{d}}y,\quad\text{as}~\delta\to 0.\end{split}$ (8.18)
We now determine $\vartheta$ and then the second moment computation is completed. Note that
${\mathcal{R}}_{\delta}=\idotsint_{{\mathbbm{R}}^{4}}\rho(x^{\prime})\rho(y^{\prime})\rho(x)\rho(y){\mathrm{d}}x^{\prime}{\mathrm{d}}y^{\prime}{\mathrm{d}}x{\mathrm{d}}y\int_{0}^{\delta^{-2}}g_{t}(x-x^{\prime})g_{t}(y-y^{\prime}){\mathrm{d}}t.$ (8.19)
We have that by the end of [14, Section 8.2],
@@
$\begin{split}R_{\delta}=\frac{1}{2\pi}\Big(&\log\delta^{-2}+\log 2-\gamma+\\ &\idotsint_{{\mathbbm{R}}^{4}}\!\rho(x^{\prime})\rho(y^{\prime})\log\!\Big[\frac{1}{(x-x^{\prime})^{2}+(y-y^{\prime})^{2}}\Big]\rho(x)\rho(y){\mathrm{d}}x^{\prime}{\mathrm{d}}y^{\prime}{\mathrm{d}}x{\mathrm{d}}y+o(1)\Big).\end{split}$
@@
$\vartheta=\log 2-\gamma+\idotsint_{{\mathbbm{R}}^{4}}\!\!\rho(x^{\prime})\rho(y^{\prime})\log\Big[\frac{1}{(x-x^{\prime})^{2}+(y-y^{\prime})^{2}}\Big]\rho(x)\rho(y){\mathrm{d}}x^{\prime}{\mathrm{d}}y^{\prime}{\mathrm{d}}x{\mathrm{d}}y+\frac{\theta}{2\pi}.$ (8.20)
8.3 Approximation by the coarse-grained model
@@
We partition the time interval $0,\delta^{-2}]$ into mesoscopic intervals
${\mathcal{I}}_{\delta,\varepsilon}(\mathrm{i}):=[(\mathrm{i}-1)\varepsilon\delta^{-2},\varepsilon\delta^{-2}],\quad\mathrm{i}=1,\cdots,\frac{1}{\varepsilon}.$ (8.21)
@@
$\begin{split}u^{\delta}[f]=&\int_{\mathbbm{R}}f(x){\mathrm{d}}x+\delta\sum\limits_{k=1}^{1/\varepsilon}\sum\limits_{1\leq\mathrm{i}_{1}<\cdots<\mathrm{i}_{k}<\frac{1}{\varepsilon}}\int_{{\mathbbm{R}}}f(\delta x_{0}){\mathrm{d}}x_{0}\sum\limits_{j_{0}:=0<j_{1}<\cdots<j_{k}}\times\\ &\bigg(\prod\limits_{\ell=1}^{k}\beta_{\delta}^{j_{\ell}-j_{\ell-1}}\idotsint\limits_{\quad t_{j_{\ell-1}+1}<\cdots<t_{j_{\ell}}\in{\mathcal{I}}_{\delta,\varepsilon}(\mathrm{i}_{\ell})\atop x_{j_{\ell-1}+1},\cdots,x_{j_{\ell}}\in{\mathbbm{R}}}\prod\limits_{i=j_{\ell-1}+1}^{j_{\ell}}g_{t_{i}-t_{i-1}}(x_{i}-x_{i-1})\rho(x_{i}){\mathrm{d}}x_{i}{\mathrm{d}}W_{t_{i}}\bigg).\end{split}$ (8.22)
By the independent increments of $W_{t}$, when computing the second moment of (8.22), the cross terms related to different $\mathrm{i}_{1}<\cdots<\mathrm{i}_{k}$ and $\mathrm{i}^{\prime}_{1}<\cdots<\mathrm{i}^{\prime}_{l}$ have no contributions. By the second moment computation in Subsection 8.2, we have an continuous analogue of Lemma 5.2, that is, as $\delta\to 0$, for fixed ${\mathbf{I}}=(\mathrm{i}_{1},\cdots,\mathrm{i}_{k})$, the related second moment converges to ${\mathcal{I}}_{\varepsilon}^{f}({\mathbf{I}})$ defined in (5.5). Thus, we can restrict the summation for $\mathrm{i}_{1},\cdots,\mathrm{i}_{k}$ to $\vec{\mathbf{{\mathcal{A}}}}_{\varepsilon}^{\text{\rm(no triple)}}$ defined in (2.42) as we have done in Section 5.
@@
$\Theta(\vec{\mathrm{i}}):=\begin{cases}\displaystyle\begin{split}\frac{\delta}{\sqrt{\varepsilon}}&\sum\limits_{k=2}^{\infty}\beta_{\delta}^{k}\int_{{\mathbbm{R}}}\rho(x_{1}){\mathrm{d}}x_{1}\times\\ &\idotsint\limits_{t_{1}<\cdots<t_{k}\in{\mathcal{I}}_{\delta,\varepsilon}(\mathrm{i})\atop x_{2},\cdots,x_{k}\in{\mathbbm{R}}}{\mathrm{d}}W_{t_{1}}\prod\limits_{i=2}^{k}g_{t_{i}-t_{i-1}}(x_{i}-x_{i-1})\rho(x_{i}){\mathrm{d}}x_{i}{\mathrm{d}}W_{t_{i}}\end{split},&\text{if}~|\vec{\mathrm{i}}|=1,\\[55.0pt] \displaystyle\begin{split}\frac{\delta}{\sqrt{\varepsilon}}&\sum\limits_{k=1}^{\infty}\beta_{\delta}^{k}\int_{{\mathbbm{R}}}\rho(x_{1}){\mathrm{d}}x_{1}\times\\ &\idotsint\limits_{t_{1}<\cdots<t_{k}\in{\mathcal{I}}_{\delta,\varepsilon}(\mathrm{i})\atop x_{2},\cdots,x_{k}\in{\mathbbm{R}}}{\mathrm{d}}W_{t_{1}}\prod\limits_{i=2}^{k}g_{t_{i}-t_{i-1}}(x_{i}-x_{i-1})\rho(x_{i}){\mathrm{d}}x_{i}{\mathrm{d}}W_{t_{i}}\times\\ &\sum\limits_{j=1}^{\infty}\beta_{\delta}^{j}\int_{{\mathcal{I}}_{\delta,\varepsilon}(\mathrm{i}^{\prime})}\int_{\mathbbm{R}}g_{s_{1}-t_{k}}(y_{1}-x_{k})\rho(y_{1}){\mathrm{d}}y_{1}{\mathrm{d}}W_{s_{1}}\times\\ &\idotsint\limits_{s_{2}<\cdots<s_{j}\in{\mathcal{I}}_{\delta,\varepsilon}(\mathrm{i}^{\prime})\atop y_{2},\cdots,y_{j}\in{\mathbbm{R}}}\prod\limits_{i=2}^{j}g_{s_{i}-s_{i-1}}(y_{i}-y_{i-1})\rho(y_{i}){\mathrm{d}}y_{i}{\mathrm{d}}W_{s_{i}}\end{split},&\text{if}~|\vec{\mathrm{i}}|>1,\end{cases}$ (8.23)